2+2=5 prove

2+2=5 prove

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

It can be done by many analytical methods.
Lets start with
-20 = -20
16-36 = 25-45
(2+2)^{2}-(2+2).9 = (5)^{2}-5.9
(2+2)^{2}-2.(2+2).\frac{9}{2} = (5)^{2}-2.5.\frac{9}{2}
(2+2)^{2}-2.(2+2).\frac{9}{2} + (\frac{9}{2})^{2} = (5)^{2}-2.5.\frac{9}{2}+(\frac{9}{2})^{2}
(2+2-\frac{9}{2})^{2} = (5-\frac{9}{2})^{2}
2+2-\frac{9}{2}= 5-\frac{9}{2}
2+2= 5
Hence, Proved.